# CAMEF: Causal-Augmented Multi-Modality Event-Driven Financial Forecasting by Integrating Time Series Patterns and Salient Macroeconomic Announcements

### v1

@@ A.1. Summarization Prompt Template
The prompts are dynamically generated based on the input. Due to the length of the event script, we first ask LLaMA-3 to summarize each chunk individually, followed by a second prompt template to generate the full summary. Below is the prompt template for summarizing individual chunks, which instructs the model to produce a summary within a specified word limit. The text within “ ” indicates the input variables:
@@ A.2. Sentiment Analytical Prompt Template
For sentiment analysis, we use a prompt that instructs the model to analyze the sentiment of a given text and rate it on a scale from 0 to 10, from extremely negative to extremely positive. The prompt also asks for an explanation of the rating. Below is the sentiment analysis prompt, where the text within “{ }” indicates the input variables:
Sentiment Analysis Prompt Please analyze the sentiment of the following {text_type} summary and rate it on a scale from 0 to 10, where: 0 = Extremely Negative 1 = Strongly Negative 2 = Very Negative 3 = Moderate Negative 4 = Slightly Negative 5 = Neutral 6 = Slightly Positive 7 = Moderate Positive 8 = Very Positive 9 = Strongly Positive 10 = Extremely Positive {text_type} summary: {text} Output the sentiment analysis as: Sentiment rating: (0 to 10), Explanation:
@@ Appendix B Implementation Details of Dataset Collection and Preprocessing
B.1. Dataset Collection Implementation Details
The dataset collection process retrieves raw macroeconomic event data files from official government sources (as illustrated in Tab. 1) to construct the CAMEF dataset. The raw data files can be presented in various formats (including PDF, HTML, and TXT). To achieve this, we leaverage the following key Python libraries to crwal the data including:
  • Requests is employed to handle HTTP requests for accessing event archives;
  • Selenium to local and download the target file from HTML headers.
  • BeautifulSoup is used to parse HTML pages and extract relevant file links;
The dataset is categorized into six macroeconomic event types: FOMC minutes, CPI reports, PPI reports, unemployment insurance releases, unemployment rate reports, and GDP advance releases. All collected files are renamed and indexed based on event type and release date.
B.2. Dataset Preprocessing Procedure
The preprocessing stage transforms the collected raw files into a structured, unified text format to ensure dataset quality and consistency while adhering to a standardized output format. The preprocessing process varies depending on the file format. For PDF files, libraries such as PyPDF2 and pdfplumber are employed to extract text and tables. For HTML files, BeautifulSoup is used for depth-first traversal of the DOM (Document Object Model) to extract structured content. For TXT files, the content is directly copied without modification, as these files are already in a structured text format.
To address potential information loss during table extraction from PDF and HTML files, we developed specialized code to convert tables into a structured format where cell contents are separated by the delimiter “—”. An example of the converted table structure is shown below:
@@ B.2. Dataset Preprocessing Procedure
The preprocessing stage ensures the dataset is clean, well-structured, and machine-compatible, enabling robust analysis and effective machine learning model training. By combining efficient crawling and preprocessing techniques, this procedure produces a high-quality dataset suitable for event-driven financial modeling.
@@ Appendix C Implementation Details
The single-modality baselines (e.g., ARIMA, DLinear, AutoFormer, FEDFormer, iTransformer, and PatchTST) were adapted from the open-source Time-Series-Library and configured to align with the event-driven forecasting context. We use
In this setup, historical time-series segments prior to each event served as input to predict trends over fixed forecasting lengths of 35, 70, and 140 steps. All models were trained for 10 epochs with a batch size of 32, maintaining consistency with the proposed method’s settings.
For multi-modality baselines, we leveraged the original TEST repository and re-implemented GPT4MTS. TEST involved a two-stage training process, combining textual prototypes and time-series decoders. GPT4MTS was modified by replacing BERT with LongFormer, allowing it to better handle longer texts. Both methods were adjusted to forecast the same time-series horizons as the single-modality models, using the same epoch and batch configurations.
Our proposed CAMEF model underwent a two-stage training process. First, the MOMENT encoder was pre-trained on time-series data to capture historical patterns. Then, the full CAMEF model was fine-tuned on aligned textual and time-series data. Some pre-trained parameters, such as MOMENT and RoBERTa (except the last layer), were frozen to retain pre-trained knowledge, while the remaining components, including parts of GPT2 and fusion layers, were trainable. This design allowed CAMEF to effectively integrate both time-series and textual features, providing a robust event-driven forecasting framework.
Further details and exact training procedures are provided in the appendix.

### v2

@@ 7. Conclusion and Future Work
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China (NSFC) under Grant Nos. 62402396 and 72471197. We would like to thank the anonymous reviewers for their valuable comments and suggestions that helped improve the quality of this paper.
@@ A.1. Summarization Prompt Template
Due to the script’s length, we firstly prompt LLaMA-3 summarizes each chunk of an event script, followed by a second prompt to generate the full summary. Below is the chunk-level prompt, which instructs the model to summarize within a word limit. Text in “ ” indicates input variables.
@@ A.2. Sentiment Analytical Prompt Template
For sentiment analysis, the prompt asks the model to rate an event scirpt’s sentiment from 0 (negative) to 10 (positive) and explain the rating. Below is the prompt template, with “ ” denoting input variables.
Sentiment Analysis Prompt Please analyze the sentiment of the following {text_type} summary and rate it on a scale from 0 to 10, where: 0 = Extremely Negative; 1 = Strongly Negative; 2 = Very; Negative; 3 = Moderate Negative; 4 = Slightly Negative; 5 = Neutral; 6 = Slightly Positive; 7 = Moderate Positive; 8 = Very Positive; 9 = Strongly Positive; 10 = Extremely Positive {text_type} summary: {text} Output the sentiment analysis as: Sentiment rating: (0 to 10), Explanation:
@@ Appendix B Implementation Details of Dataset Collection and Preprocessing
B.1. Dataset Collection
To construct the CAMEF dataset, raw macroeconomic event data are crawled from official government sources (Tab. 1) in various formats (PDF, HTML, TXT). We utilize the following Python libraries:
  • Requests – to send HTTP requests for archive access;
  • Selenium – to locate and download files via HTML headers;
  • BeautifulSoup – to parse HTML and extract file links.
The dataset covers six event types: FOMC minutes, CPI, PPI, unemployment insurance, unemployment rate, and GDP advance releases.
B.2. Preprocessing
Raw files are converted into a unified text format. PyPDF2 and pdfplumber extract content from PDFs, while BeautifulSoup parses HTML using depth-first traversal. TXT files are used as-is. We convert all the tables into structured text using the delimiter “—”, as shown:
@@ B.2. Preprocessing
This preprocessing ensures clean, consistent, machine-readable data, supporting robust analysis and model training.
@@ Appendix C Implementation Details
Single-modality baselines (ARIMA, DLinear, AutoFormer, FEDFormer, iTransformer, PatchTST) were adapted from Time-Series-Library to the event-driven setting, using pre-event segments to predict 35-, 70-, and 140-step trends. All models were trained for 10 epochs with batch size 32.
For multi-modality baselines, we used TEST and a re-implemented GPT4MTS (with LongFormer replacing BERT). Both were aligned to the same forecasting horizons and training settings.
CAMEF followed a two-stage pipeline: MOMENT was pre-trained on time series, then the full model fine-tuned on aligned text–series pairs. MOMENT and RoBERTa (excluding final layer) were frozen, while GPT2 and fusion layers were trainable, enabling robust multimodal forecasting.